### v2

# Multi: Multimodal Understanding Leaderboard with Text and Images

Abstract
Rapid progress in multimodal large language models (MLLMs) highlights the need to introduce challenging yet realistic benchmarks to the academic community, while existing benchmarks primarily focus on understanding simple natural images and short context. In this paper, we present Multi as a cutting-edge benchmark for evaluating MLLMs on understanding complex tables and images and reasoning with long context. Multi provides multimodal inputs and requires responses that are either precise or open-ended, reflecting real-life examination styles. Multi includes over 18,000 questions, and challenges MLLMs with a variety of tasks, ranging from formula derivation to image detail analysis and cross-modality reasoning. We also introduce Multi-Elite, a 500-question selected hard subset, and Multi-Extend with more than 4,500 external knowledge context pieces. Our evaluation indicates significant potential for MLLM advancement, with GPT-4V achieving a 63.7% accuracy rate on Multi, in contrast to other MLLMs scoring between 28.5% and 55.3%. Multi serves not only as a robust evaluation platform but also paves the way for the development of expert-level AI. Details and access are available at: https://OpenDFM.github.io/MULTI-Benchmark/.
Multi: Multimodal Understanding Leaderboard with Text and Images
Zichen Zhu and Yang Xu and Lu Chen22footnotemark: 2and Jingkai Yang and Yichuan Ma Yiming Sun and Hailin Wen and Jiaqi Liu and Jinyu Cai Yingzi Ma and Situo Zhang and Zihan Zhao and Liangtai Sun and Kai Yu22footnotemark: 2 X-LANCE Lab, Department of Computer Science and Engineering MoE Key Lab of Artificial Intelligence, SJTU AI Institute Shanghai Jiao Tong University, Shanghai, China {JamesZhutheThird, xuyang0112, chenlusz, kai.yu}@sjtu.edu.cn
Figure 1: An example of Multi. English translations of Chinese text are shown for better readability. The markdown format remains as it is.
Figure 2: The overview of Multi.
1 Introduction
The rapid advancement in large-scale language models (LLMs) has led to significant achievements in natural language processing and related disciplines. Yet, human communication and understanding extend beyond language, encompassing images, tables, mathematical and chemical formulas, graphs, diagrams, cartoons, posters, and other visual mediums. They play a crucial role in conveying information, particularly in scientific areas. Therefore, there’s a growing interest in developing Multimodal LLMs (MLLMs) capable of processing and generating across various modalities, including visual ones, and performing tasks that require cross-modal reasoning.
Evaluating MLLMs presents unique challenges. Current benchmarks Lu et al. (2022); Li et al. (2023b); Yue et al. (2023) either focus narrowly on natural scene images or are simplistic, failing to thoroughly assess the models’ abilities. Many scientific benchmarks Sun et al. (2023a); Huang et al. (2023) rely on multiple-choice questions with a single answer, which may not accurately gauge a model’s comprehension and can lead to superficial learning, i.e., the model will not look into other choices if the correct choice is straightforward. A more robust, detailed, and multi-scale dataset is necessary to effectively evaluate MLLMs under diverse conditions and scenarios. Current benchmarks mentioned above are evaluated with English context, while the rapid progression of Chinese MLLMs highlights the need for a Chinese multimodal benchmark with Chinese contents both in text and image and brings new challenges to the community.
In this paper, we introduce Multi, a novel benchmark named Multimodal Understanding Leaderboard with Text and Images, specifically designed to evaluate multimodal LLMs on cross-modal questions. Multi comprises 18,430 questions sourced from various educational and online materials, with most questions undergoing multiple rounds of human annotation for quality assurance. These questions cover a variety of scientific disciplines, including mathematics, physics, computer science, etc., and also pose significant challenges to intricate image reasoning. Multi serves as the first benchmark incorporating driving tests and administrative aptitude tests in China. The questions are crafted to test understanding and generation in various formats and complexity levels and are categorized into multiple-choice (with single or multiple correct answers), fill-in-the-blank, and open-ended questions.
To further challenge multimodal LLMs, we develop two subsets within Multi: Multi-Elite  consists of 500 carefully selected tough questions aiming to probe the limits of these models, and Multi-Extend  featuring 4,596 knowledge pieces tests the models’ capabilities of learning and knowledge transfer. These subsets offer deeper insights into the strengths and weaknesses of multimodal LLMs, fostering new research avenues. An example of Multi is shown in Figure 1, and more are presented in Appendix F.
We conduct comprehensive experiments on Multi using leading-edge multimodal and single-modality LLMs. Our findings reveal that multimodal LLMs still lag behind human performance in many aspects of Multi, highlighting challenges like cross-modal alignment, logical reasoning, mathematical computations, and image comprehension. Results show that the benchmark is challenging for current models, not to mention the Multi-Elite set where GPT-4V only gets a 14.0% score, and most of the other models get a score near random, indicating a large space for improvement.
In conclusion, We make the following contributions in this work:
  • We propose Multi, a substantial and challenging multimodal benchmark focusing on Chinese scientific questions, designed to evaluate multimodal LLMs.
  • We introduce Multi-Elite and Multi-Extend sets to test models’ bottleneck and in-context learning abilities, aiming for a more nuanced evaluation of multimodal LLMs.
  • We present detailed experiments with various state-of-the-art multimodal and single-modality LLMs on Multi, providing both qualitative and quantitative insights into their performance.
  • We make the Multi leaderboard, dataset, evaluation code, and the two subsets available to the research community, encouraging further participation and advancement in the field of multimodal LLMs.
2 Related Works
Multimodal Large Language Models (MLLMs).
With advancements in aligning features across multiple modalities, like CLIP Radford et al. (2021) and ALBEF Li et al. (2021), recent studies have explored projecting vision features into the latent space of LLMs, aiming to enhance their capabilities of comprehending visual information. For example, BLIP-2 Li et al. (2023c) pioneers this approach by employing Q-Former to translate image features into text representations. Following this, LLaVA Liu et al. (2023b), MiniGPT-4 Zhu et al. (2023), and InstructBLIP Dai et al. (2023) have introduced visual instruction tuning to bolster the capability of MLLMs of following instructions. Our primary focus is on the proficiency of MLLMs in comprehending instructions in Chinese, which are divided into two main branches: open-source models, which typically build upon existing Chinese LLMs or are fine-tuned on Chinese instruction datasets, examples of which include Chinese-LLaVA LinkSoul-AI (2023), VisualGLM Du et al. (2022), VisCPM Hu et al. (2023), Qwen-VL Bai et al. (2023a), InternVL Zhang et al. (2023a), Yi-VL 01.ai (2023); and closed-source models, which are often highly powerful, multi-lingual systems such as GPT-4V(ision) OpenAI (2023b) and Gemini Team (2023). In this paper, we intend to evaluate these models across a range of scientific fields on the Multi benchmark, offering an extensive assessment and guidance for the onward trajectory of Chinese MLLMs.
Benchmarks for MLLMs.
In assessing MLLMs, traditional methods primarily rely on established vision-language (VL) benchmark datasets. Renowned benchmarks such as VQA Goyal et al. (2017), OK-VQA Antol et al. (2015), GQA Hudson and Manning (2019), and MSCOCO Lin et al. (2014) are tailored to specific VL tasks like image captioning, open-domain visual question answering, and visual reasoning. While the evaluation based on standard benchmark datasets yields significant insights into MLLMs’ capabilities, these approaches may not entirely capture their comprehensive intelligence in real-world scenarios. Therefore, a diverse array of benchmarks has been developed to examine MLLMs on dealing with various tasks in real world. Benchmarks like LLaVA-Bench Liu et al. (2023b), MMBench Liu et al. (2023c), MM-VET Yu et al. (2023), TouchStone Bai et al. (2023b), MLLM-bench Ge et al. (2023), and SEED-BenchLi et al. (2023b, a), for instance, leverage GPT to evaluate the relevance and helpfulness of human-like long responses in the reality. POPE Li et al. (2023d) and HallusionBench Liu et al. (2023a) introduce various analytical criteria for the holistic evaluation of MLLMs’ hallucinations. Furthermore, M3Exam Zhang et al. (2023b), SciGraphQA Li and Tajbakhsh (2023), MathVista Lu et al. (2023), AGIEval Zhong et al. (2023), and MMMU Yue et al. (2023) consider MLLMs as experts to extend the evaluation scope by incorporating advanced perception and reasoning within domain-specific knowledge, for example, scientific questions and driving tests. The works most related to us are M3Exam, ScienceQA, SciEval Sun et al. (2023a) and C-Eval Huang et al. (2023). Our approach distinguishes itself by offering a broader spectrum of question types compared to the first two and supports a multimodal evaluation in contrast to the last two.
3 The Multi Benchmark
We propose Multi, a Multimodal Understanding Leaderboard with Text and Images, which can serve as a challenging and diverse benchmark for the MLLM community. The detailed statistics are provided in Appendix A.
3.1 Data Construction Process
Figure 3: The construction pipeline of Multi.
The data construction pipeline is shown in Figure 3. To develop Multi, we follow several key steps to ensure high-quality and precise annotation. Firstly, we crawl open-source raw question data from the Internet and transcript close-source exams from paper documents. Secondly, we format each question and knowledge piece into markdown and  formula format to maintain precision and quality. Thirdly, we revise and refine each question multiple times to prevent data leakage and increase difficulty. Lastly, We rate every question based on its difficulty and content richness.
Data Source
We collect more than 2.7M raw data from the Internet, ranging from exams and quizzes from Chinese junior and senior schools and several society exams. We design an algorithm to pick out a proportion of the questions as the fundamental data of our benchmark. The selection is based on the questions’ text length, number of images, corresponding subjects, and knowledge pieces, to reach a higher diversity of questions and coverage of knowledge. The details are presented in Appendix D. We also collect questions from internal exams and practices of several top universities. After the selection, we obtain over 18K questions as the raw data.
Data Process and Annotation
The data process and annotation for our dataset involve a comprehensive series of steps to ensure high-quality, diverse content.
In the Data Pre-process stage, raw data with formats like HTML, photocopy, hand script, or plain text are refined by removing irrelevant HTML tags, converting text styles into markdown format, and transcribing math functions and chemical structures into  format, with complex tables saved as screenshot images after HTML rendering. OCR tools are utilized for text conversion from photocopies and hand scripts.
During the Data Annotation stage, an online platform facilitates annotators, primarily skilled undergraduates (involved in the work as authors), in tasks across format, content, label, and semantic levels. This includes converting content into markdown and , splitting sub-questions into individual ones, evaluating the difficulty and quality, and correcting errors for factual accuracy.
The Data Post-process stage employs strategies like formation, disambiguration, distillation, and transformation to enhance question difficulty and diversity, including modifying question formats and reducing assistance information.
Throughout these stages, we process 2.7 million questions in total and pick out 18,430, incorporating 23,320 scoring points, 7,658 images, and 4,595 knowledge pieces. Multi highlights a broad diversity in question types, including multiple-choice questions with both single and multiple answers, along with fill-in-the-blank and open-ended writing questions enriching the testing scenarios.  [^1] For the sake of simplifying writing, in the following paragraphs we may use abbreviations. We denote multiple-choice questions with a single answer as SA or Single Answer Choosing and those with multiple answers as MA or Multi Answer Choosing. We use FB for fill-in-the-blank questions and OP for open-ended writing questions. The stages during data processing and annotation significantly increase the diversity and difficulty of the dataset. For details of data processing and annotation, please refer to Appendix E.
3.2 The Multi-Elite Set
We select an additional set of 500 questions to create the advanced dataset. This set is comprised of objective questions, i.e. multiple-choice and fill-in-the-blank questions. The questions are averagely distributed in all of the subjects and education levels, evaluated as with high difficulty and quality by annotators, and with rich text and image content. The evaluation results presented in § 4 are also referred to in the selection, where the results of GPT-4VOpenAI (2023b) are given the most consideration.
3.3 The Multi-Extend Background Knowledge Dataset
External knowledge is crucial to provide critical information that assists in solving questions using the In-Context Learning (ICL) abilities. Some of the raw questions retrieved from the Internet have corresponding knowledge pieces attached. We also collect more knowledge pieces for uncovered questions with the assistance of LLMs and outer knowledge source (e.g. New Bing [^2] https://bing.com/new and Wikipedia [^3] https://wikipedia.org). We conduct annotations on these knowledge pieces to confirm the correctness of the content and present them in the Multi-Extend dataset. This dataset consists of about 4.6K knowledge pieces, designed to test the in-context learning abilities and knowledge transfer skills of models. This dataset provides comprehensive insights into the capabilities and limitations of multimodal LLMs, opening new pathways for research exploration.
3.4 Comparison with Existing Benchmarks
Benchmark Lang Size Image Answer Type Source
Sub Q Ana Img Kn NI SI MI SA MA FB OP
VQA Antol et al. (2015) en 36 764K - 265K - Repurposed
ScienceQA Lu et al. (2022) en 21 21K 19K 10K 0.3K Textbooks
SciBench Wang et al. (2023) en 6 0.8K - 0.1K - Textbooks
M3Exam Zhang et al. (2023b) 9 langs 4 12K - 3.1K - Exams
AGIEval Zhong et al. (2023) zh, en 20 8K a few - - Exams
MMBench Liu et al. (2023c) en 20 3K - 3K - Web, Repurposed
SEED-Bench Li et al. (2023b) en 12 19K - 19K+ - Anno.
SEED-Bench-2 Li et al. (2023a) en 27 24K - 22K+ - Anno.
MLLM-Bench Ge et al. (2023) en 42 0.4K - 0.4K - Anno.
Touchstone Bai et al. (2023b) en 27 0.9K - 0.9K - Anno.
C-Eval Huang et al. (2023) zh 52 14K a few - - Exams, Web
SciEval Sun et al. (2023a) en 3 18K - - - Web, Repurposed
MMMU Yue et al. (2023) en 30 12K 2K 11K+ - Anno., Web, Textbooks
Multi (ours) zh 23 18K 10K+ 7.7K 4.6K Anno., Exams, Web
Table 1: The comparison between Multi and other existing benchmarks. Sub: Subject or Field, Q: Question, Ana: Analysis or Explanations, Img: Images, Kn: Knowledge or Lecture. NI: the question with pure text, SI: the question with a single image, MI: the question with multiple images. SA: multiple-choice question with single correct answer, MA: multiple-choice question with multiple correct answers, FB: fill-in-the-blank question (no more than 10 words), OP: open-ended writing question (more than 10 words). Anno.: Annotation
Multi demonstrates a comprehensive blend of features that surpasses existing benchmarks in several dimensions. Notably, Multi covers a wide array of subjects and a substantial number of questions (18K), as well as over 10K analysis and 4.6K extensive knowledge content, which is considerably larger than most benchmarks, ensuring a broad and diverse testing environment. Multi possesses 7.7K images, which is essential for benchmarking MLLMs that require visual understanding alongside textual information. The inclusion of both single and multiple image questions, as well as a variety of answer types, makes Multi a versatile and challenging benchmark. Furthermore, the questions without images also test the MLLMs’ ability on dealing with plain text information. Meanwhile, the various sources, complex annotation, and processing stages provide sufficient augmentation to alleviate data leakage. Multi not only encompasses variations of classic questions but also includes recently updated questions, which significantly enhances its diversity.
We list the features of existing benchmarks and make a comparison with Multi in Table 1. We believe that Multi assembles the most advantages of the existing benchmarks and is sure to provide a good option for the community to test the capabilities of their Vision LLMs.
4 Experiments
4.1 Models
We evaluate a wide range of MLLMs that support Chinese, including Chinese-LLaVA LinkSoul-AI (2023), Qwen-VL Bai et al. (2023a), VisCPM Hu et al. (2023), VisualGLM Du et al. (2022), InternVL Chen et al. (2023), Yi-VL 01.ai (2023), Gemini Vision Team (2023), and GPT-4V OpenAI (2023b). We evaluate these models with both multimodal input and text-only input to verify the information gain of input images. We also select several most capable LLMs for comparison with text-only input, including DFM-2 Chen et al. (2022), MOSS Sun et al. (2023b), ChatGPT OpenAI (2022), GPT-4 OpenAI (2023a), and Gemini Team (2023), and the performance of these models on questions with images will reflect their abilities on finding loss of information. Model specifications are listed in Table 10. Due to the API request rate limit of Gemini and GPTs, ablation studies are mostly performed on weight-accessible models. We choose the checkpoints with the largest model size and the latest version, and use FP16 or INT4 quantization to accelerate inference if officially provided. We follow the official guidelines to prompt each model so that the outputs go in the desired way.
4.2 Settings
Prompt
We use specialized prompts for each question, an example shown in Figure 4. The prompts are designed carefully according to the features of each type of question and the answer patterns expected. We also modify the input format to fit into official inference guidelines. The complete collection of prompts is presented in Appendix C.
Figure 4: An example of the prompts used when evaluating a multiple-choice question with image context, knowledge piece and single correct answer.
Image
Multi includes questions with either none, single, or multiple images. Most MLLMs accept text accompanied by one image as input or a pure-text input. For questions with a single image, the image and text are fed in one turn. We simply drop image information when evaluating LLMs.
For pure-text questions, we use the text as input. For some models like VisCPM, InternVL and Yi-VL which compulsorily demand an image in each turn, we feed the model a blank image with color set to RGB(0,0,0) along with plain text in evaluation. For efficiency, results of GPT-4 and Gemini on pure-text questions are directly used as the results of GPT-4V and Gemini Vision respectively.
[…]
4.3 Metrics
We focus on objective questions with a certain answer, including multiple-choice and blank-filling questions. We also give a score to each subjective open-ended question based on the similarity to the reference answer. The metrics we use for each type of question:
Multiple-choice with Single Answer (SA)
Each question worth one point. We calculate the accuracy of the given answer.
Multiple-choice with Multiple Answers (MA)
We define the total points of an MA question as the number of correct choices, and each correct choice selected is rewarded one point. If the given answer contains any wrong choice, the score will be counted to zero. We report the score ratio (# points / # total points) as the metric. We also report accuracy as a more rigorous metric, where only correctly giving all the choices without wrong ones will be granted points.  [^4] For example, a question with correct answer ACE worth 3 points, and answer AC will be granted 2 points and answer BC or ABCE will be granted 0 points. However, on calculating accuracy none will be counted, and only ACE will be calculated as correct.
Fill in the Blank (FB)
We define the total points of a blank-filling question as the number of the blanks marked as [MASK]. It is required in prompts that each line of the given answer corresponds to a blank in order. We follow the most strict standard of exact match. Therefore, only answers exactly matching the standard answers will be granted points. We report the score ratio as the final metric.
Open-ended Question (OP)
The points and counting method is similar to FB, but we use a loose standard and report normalized ROUGE-L Lin (2004) score for each point. Please note that the reference answer may be concise or in detail, and there could be other possible answers.
4.4 Main Experiment Results
We report the overall and field-specific performance of tested models on the whole benchmark in Table 2, 3, and 4.
Model Overall NI SI MI
Puretext (LLM)
MOSS 32.6 36.1 27.3 17.1
DFM-2.0 49.7 63.0 28.7 11.3
Gemini 52.2 62.5 36.2 18.3
ChatGPT 35.9 54.0 6.8 5.1
GPT-4 50.2 74.5 11.3 8.8
Text+Image (MLLM)
Chinese-LLaVA 28.5 32.3 22.6 17.8
VisualGLM 31.1 35.1 25.2 9.7
VisCPM 33.4 36.8 28.4 16.6
Qwen-VL 39.0 43.2 32.7 20.7
InternVL 44.9 50.9 35.5 25.1
Yi-VL 55.3 63.8 42.0 24.5
Gemini Vision 53.7 62.5 40.0 24.5
GPT-4V 63.7 74.5 46.9 28.1
Table 2: The main performance of models evaluated on Multi. NI: the question with no image, SI: the question with a single image, MI: the question with multiple images.
Overall comparison.
We report the overall performance in Table 2. The most powerful competitor, GPT-4V, achieves a mere 63.7% score, underscoring the benchmark’s complexity and challenge. Yi-VL outperforms other open-source models, but there still remains a notable gap with GPT-4V, and those smaller models do not get as much as half of the scores.
Comparison by number of images.
In Table 2, we also present the performance categorized by image number. For MLLMs, a higher score on the Non-Image (NI) set suggests improved performance on multimodal questions, including the Single Image (SI) set and Multiple Image (MI) set. It is evident that questions requiring more images are more challenging. A significant drop in performance is observed when answering questions with more than one image. Only GPT-4V (28.1%) manages to exceed the average baseline set by random guessing.
Conversely, for LLMs, there exists a reverse correlation between scores on the NI set and those on the SI and MI sets. This is because we prompt the model to determine whether visual information is necessary for answering a question and if so the model needs to refuse to answer. Less capable models may simply make a guess, but more sophisticated models tend to withhold an answer, resulting in lower but more reliable overall scores. The results on SI and MI sets for LLMs indicate a long way before mitigating hallucination.
Model SA MA MA Acc. FB OP
Puretext (LLM)
MOSS 38.5 33.1 6.8 2.7 8.7
DFM-2.0 55.8 53.9 29.7 13.3 10.3
Gemini 58.2 52.7 22.8 29.1 7.9
ChatGPT 40.0 39.4 17.9 10.5 7.7
GPT-4 51.3 60.0 53.1 32.9 6.8
Text+Image (MLLM)
Chinese-LLaVA 34.5 26.9 3.9 2.4 8.4
VisualGLM 37.9 30.2 1.9 0.7 3.6
VisCPM 41.7 27.7 0.0 3.8 14.1
Qwen-VL 49.8 29.4 2.8 5.8 13.7
InternVL 56.4 33.4 2.1 14.2 13.1
Yi-VL 61.3 42.0 36.4 14.6 8.9
Gemini Vision 59.4 54.4 24.3 30.5 12.5
GPT-4V 67.1 70.6 58.2 42.4 11.7
Table 3: Performance of models on each type of questions of Multi. MA Acc.: Accuracy of MA questions.
Comparison by question type.
In Table 3, we present the performance categorized by question type. A majority of the models achieve their highest scores on the Single Answer Choosing (SA) set, with lower performance on the Multiple Answers Choosing (MA) set. A notable discrepancy is observed between the scores for the MA set and its accuracy, highlighting the smaller models’ inability to identify all correct options accurately.
For the Fill-in-the-Blank (FB) set, which requires short but exact matches, the scores further decline. This is partially due to failure to follow the specified instructions, often leading to correct responses being presented in an unacceptable format.
Furthermore, we note significantly lower scores on the Open-ended Writing (OP) set in comparison to the FB set. VisCPM stands out but only with the best score of 14.1% on the OP set, suggesting that our dataset minimizes the risk of data leakage and poses considerable challenges for models in generation across modalities.
Model JuH SeH Uni Driv AAT
Puretext (LLM)
MOSS 21.2 26.7 23.8 44.1 25.5
DFM-2.0 42.3 42.5 35.7 66.3 3.9
Gemini 47.7 42.3 41.4 66.9 22.5
ChatGPT 31.6 23.7 34.9 52.1 1.3
GPT-4 49.2 33.7 55.1 69.9 0.9
Text+Image (MLLM)
Chinese-LLaVA 21.1 25.4 20.7 35.8 21.8
VisualGLM 22.2 25.6 23.6 40.9 24.9
VisCPM 25.2 28.1 23.0 43.4 23.7
Qwen-VL 32.6 32.9 27.2 49.3 26.4
InternVL 39.3 36.5 30.6 57.7 24.8
Yi-VL 46.6 46.0 45.4 71.1 26.5
Gemini Vision 48.2 45.2 41.7 67.4 27.0
GPT-4V 58.5 52.9 59.0 80.1 26.2
Table 4: Performance of models on each subject of Multi. JuH: level of Junior High school, SeH: level of Senior High school, Uni: level of University, Driv: Chinese driving test, AAT: Administrative Aptitude Test.
Comparison by education level and subjects.
In Table 4, we present the performance categorized by educational levels and subjects. The performance trends for high school and university level questions remain consistent with the overall results observed. For questions at the society level, we anticipate higher scores on the Driving Test. This may be caused by a larger percentage of judgmental questions (in the format of SA with two options), as well as its nature with knowledge of regulations.
Furthermore, questions from the Administrative Aptitude Test (AAT), which typically include at least one image and often examine skills on image pattern recognition (illustrated in the first two examples in the left column of Figure 9), tend to have scores around or below randomly choosing baseline. Even the strongest competitor, GPT-4V, shows limited success, with a performance of only 27.0% on these questions as detailed in the study cited in the paper OpenAI (2023b). This underscores the significant challenge posed by multimodal questions. Notably, the stronger LLMs, specifically DFM-2.0 and the text-only versions of GPT perform poorly on AAT questions as expected, as they often reject answering the majority of them.
4.5 Ablation Study on Image Information Gain
Model NI SI MI
w/o. image w. caption w. ocr w. image w/o. image w. caption w. ocr w. image
Puretext (LLM)
MOSS 36.1 27.3 27.3 (+0.0) 27.6 (+0.3) - 17.1 20.7 (+3.6) 19.0 (+1.9) -
ChatGPT 54.0 6.8 9.9 (+3.1) 6.6 (-0.2) - 5.1 10.7 (+5.6) 5.5 (+0.4) -
DFM-2.0 63.0 28.7 30.2 (+1.5) 33.4 (+4.7) - 11.3 15.6 (+4.3) 14.9 (+3.6) -
Text+Image (MLLM)
Chinese-LLaVA 32.3 26.1 26.3 (+0.2) 25.5 (-0.6) 22.6 (-3.5) 17.6 19.9 (+2.3) 19.6 (+2.0) 17.8 (+0.2)
VisualGLM 35.1 20.8 21.4 (+0.6) 20.4 (-0.4) 25.2 (+4.4) 15.3 15.1 (-0.2) 14.5 (-0.8) 9.7 (-5.6)
VisCPM 36.8 27.1 27.6 (+0.5) 27.2 (+0.1) 28.4 (+1.3) 24.8 21.6 (-3.2) 20.9 (-3.9) 16.6 (-8.2)
Qwen-VL 43.2 30.7 30.3 (-0.4) 31.0 (+0.3) 32.7 (+2.0) 25.5 25.0 (-0.5) 26.2 (+0.7) 20.7 (-4.8)
InternVL 50.9 33.4 33.3 (-0.1) 33.1 (-0.3) 35.5 (+2.1) 24.8 21.9 (-2.9) 22.9 (-1.9) 25.1 (+0.3)
Gemini/Vision 62.5 36.2 36.9 (+0.7) 38.4 (+2.2) 40.0 (+3.8) 18.3 23.2 (+4.9) 18.6 (+0.3) 24.5 (+6.2)
Yi-VL 63.8 39.9 38.7 (-1.2) 39.4 (-0.5) 42.0 (+2.1) 24.1 26.5 (+2.4) 24.2(+0.1) 24.5 (+0.4)
GPT-4/V 74.5 11.3 9.7 (-1.6) 1.9 (-9.4) 46.9 (+35.6) 8.8 9.4 (+0.6) 3.1 (-5.7) 28.1 (+19.3)
average +0.30 -0.35 +5.98 +1.54 -0.30 +0.98
Table 5: Performance of models evaluated on the image set of Multi.
To assess the necessity of images in Multi for solving problems, we conduct an ablation study where we either remove images from the SI and MI sets or substitute them with textual descriptions, such as captions and OCR-derived text. We utilize BLIP2 Li et al. (2023c) for generating image captions and EasyOCR [^5] https://pypi.org/project/easyocr/ to extract text from images. The results are shown in Table 5.
For questions that incorporate a single image (as indicated in the SI column), the presence of images significantly aids in answering the questions, with an average performance boost of 5.98%. Notably, GPT-4V experiences a substantial increase of 35.6% in performance, primarily due to its tendency to abstain from answering in the absence of images.
In settings where images are omitted and replaced by their textual descriptions (captions or OCR text), there’s a marginal improvement of 0.30% observed with captions, but a minor reduction of -0.35% with OCR text. Captions, which generally summarize the images, introduce bilingual elements to the models and usually miss details. OCR text, while detailed, lacks spatial information and is not universally applicable, as some images contain no text at all. Both forms of textual information lower the models’ refusal rate, and LLMs benefit more from these than MLLMs. However, they potentially complicate reasoning processes. Nevertheless, a generic caption is found to be more beneficial than scattered OCR fragments.
For questions that involve multiple images (as discussed in the MI column), we categorize models into three groups: 1) Close-source models, specifically GPT-4V and Gemini Pro, which leverage all images and achieve significant improvement. 2) Open-source models capable of handling multiple images within a dialogue or at a time, namely VisualGLM, VisCPM, and Qwen-VL, all of which exhibit a notable performance decline. 3) Open-source models without multi-image support, like Chinese-LLaVA, InternVL, and Yi-VL, show slight improvements. The second group’s decline could be attributed to their inability to utilize conversation history effectively and remember previously seen images. The third group’s limitation likely stems from providing only the first image, insufficient for comprehending all necessary information to answer the question, but to some degree avoiding distraction.
4.6 Evaluation on Multi-Elite
Model Overall SA MA MA Acc. FB NI SI MI
Chinese-LLaVA 12.3 15.7 13.1 1.0 1.6 13.7 11.0 15.3
VisualGLM 12.8 14.5 16.6 0.0 0.8 16.2 11.7 6.8
VisCPM 13.0 10.4 22.0 0.0 0.8 10.3 14.2 15.3
Qwen-VL 10.5 7.2 19.3 1.9 0.8 8.5 10.8 16.9
InternVL 20.7 24.8 23.2 0.0 4.8 17.9 21.0 28.8
Yi-VL 26.2 33.0 29.0 8.7 3.2 32.5 22.7 25.4
Gemini Vision 12.4 5.3 21.2 5.8 12.0 6.8 12.0 37.3
GPT-4V 14.0 5.3 25.5 15.4 12.0 7.3 14.9 33.9
Table 6: Performance of models on Multi-Elite.
We conduct evaluations on Multi-Elite, as outlined in Table 6, which includes 500 specifically chosen questions. These questions are selected based on pre-annotated quality and difficulty scores, in addition to the evaluation results on Multi discussed in § 4.4. The selection aims to ensure a distribution that mirrors Multi’s but also brings challenge to strong MLLMs. Yi-VL achieves the highest score on Multi-Elite with 26.2%, while scores for other models vary between 10.5% and 20.7%. This highlights the substantial challenge presented by Multi-Elite, indicating significant potential for improvement in tackling extremely difficult questions that require in-depth image understanding and intricate reasoning across modalities. It is important to highlight the accuracy of multiple answers choosing (MA Acc.) as the most demanding task for MLLMs, necessitating a thorough grasp of the relationships between the choices and the questions, and reflecting model reliability of selecting all answers correctly.
4.7 Evaluation on Multi-Elite with Multi-Extend
Model window size w/o. kn w. kn
InternVL 768 tokens 20.7 19.9 (-0.8)
Yi-VL 4,096 tokens 26.2 21.4 (-4.8)
Qwen-VL 8,192 tokens 10.5 13.0 (+2.5)
Gemini Vision 30,720 tokens 12.4 17.0 (+4.6)
GPT-4V 128,000 tokens 14.0 21.3 (+7.3)
Table 7: Performance of models with Multi-Extend on Multi-Elite.
The significant challenges posed by Multi-Elite prompt further investigation into the In-Context Learning (ICL) capabilities of MLLMs through the utilization of the Multi-Extend knowledge set. This set is designed to include relevant concepts and frequently utilized solutions related to the problems. The study is conducted on several MLLMs, with the prompts for incorporating these knowledge pieces shown in Figure 5, and the results are listed in Table 7. Notably, the average number of tokens per question escalates from 65 to 250, and further to 850, following the integration of prompts and the adoption of Multi-Extend, with the most extensive examples surpassing 10,000 tokens. Multi-Extend poses a significant challenge in terms of the necessary window size and the capacity to handle lengthy contexts. It is observed that models equipped with larger window sizes, i.e. Gemini Vision and GPT-4V, benefit more from Multi-Extend, whereas there is a notable decline in performance for MLLMs with smaller window sizes. The increase in tokens may also present a hurdle for models, as the concise question may become overshadowed by the extensive context.
4.8 Takeaways
  • GPT-4V demonstrates the highest performance with a 63.7% score, indicating a significant challenge of Multi, while Yi-VL leads among open-source models.
  • MLLMs show a performance drop in questions requiring more images, with only GPT-4V exceeding a basic guessing baseline in multi-image scenarios.
  • LLMs show a reverse correlation in performance between non-image and single/multiple image sets, highlighting the challenge of avoiding hallucination in visual questions.
  • Models generally perform better on questions requiring shorter answers, i.e. SA > MA > FB > OP. The results of MA Acc. emphasize the importance of balancing recall and precision.
  • Performance trends are consistent across educational levels, with lower scores on AAT questions due to their multimodal complexity.
  • The inclusion of images significantly boosts question-answering performance, with captions offering a slight improvement and OCR text potentially complicating reasoning processes.
  • In the Multi-Elite evaluation, Yi-VL achieves the highest 26.2% score, illustrating the difficulty of Multi-Elite and the need for advanced image understanding and reasoning across modalities.
  • The aid of Multi-Extend help improve performance on models with long window sizes, yet it may yield adverse effects on less capable models.
5 Conclusion
In this paper, we introduce Multi, a comprehensive and challenging benchmark designed to rigorously evaluate the performance of MLLMs in detailed cross-modality understanding and scientific reasoning. Our experiments with state-of-the-art models like Qwen-VL, InternVL, Yi-VL, Gemini, and GPT-4 demonstrate that while these models exhibit promising capabilities, there remains a significant gap compared to human performance, particularly in tasks involving cross-modal alignment, logical reasoning, and complex comprehension. This underscores the need for continuous research and development in this domain.
The creation of the Multi-Elite and Multi-Extend subsets further contributes to the field by providing insights into the strengths and limitations of current MLLMs. These subsets challenge the models’ learning and reasoning abilities and encourage the development of more sophisticated and robust multimodal understanding systems.
Multi benchmark opens new avenues for research, particularly in enhancing the MLLMs’ ability to integrate and reason over diverse data types, including images, text, and structured data. Future work may focus on expanding the benchmark to include more diverse modalities and question types, further pushing the boundaries of what MLLMs can achieve. By making Multi publicly available, we hope to foster a collaborative environment where researchers can continuously test and improve the capabilities of MLLMs, driving the field toward the development of truly intelligent and versatile AI systems.
Limitations and Future Work
Multilingual Capabilities
Multi predominantly features simplified Chinese and mainly focuses on subjects taught in Chinese schools, with limited English multimodal content that’s relatively straightforward for LLMs. We plan to include translations in future versions. Nonetheless, the presence of Chinese characters in figures poses a significant challenge for MLLMs trained on different linguistic datasets.
Use of Explanations
While we have annotated explanations in detail, the utilization in subsequent studies remains limited. These explanations could potentially serve as valuable training data for model fine-tuning and few-shot learning using methods like CoT (Chain-of-Thoughts) or RAG (Retrieval Augmented Generation) and may aid in evaluating reasoning skills.
Metrics for evaluating blank-filling, open-ended writing and others
Our evaluation primarily uses exact match, which might be overly stringent for assessing MLLMs’ true capabilities. Assessing open-ended writing tasks that require complex knowledge and reasoning is still a challenge. We also have 100 questions that do not belong to traditional categories, such as questions requiring geographic drawing, and the evaluation of them will be even more challenging. Now that only a few studies Wang et al. (2023) involve human evaluation, developing automatic and reliable methods remains an open research area.
Adaptation to various MLLMs
Although we have tested several MLLMs, numerous others exist and new ones are continuously emerging. We encourage the community to evaluate their MLLMs using our benchmark to gauge their cognitive reasoning abilities. We will test more models as soon as the multilingual version is released.
Expansion to more modalities and subjects
Our benchmark currently focuses on static images, but incorporating other modalities like audio and video, and subjects like art, music theory, medicine, and sports could present new topics. Thus, expanding our question set to cover these areas is a promising direction for future research.
Appendix A Statistics
We providedetailed statistics in Table 8. One question may contain more than one scoring points as mentioned in § 4.3.
Statistics Number Points
Total Problems 17251 -
Total Questions 18430 -
Total Points 23320 -
Total Images 7658 -
Total Knowledge 4595 -
Multiple Choices 16100(87.36%) 19904(85.35%)
 - Single Answer 13963(75.76%) 13963(59.88%)
 - Multiple Answers 2137(11.60%) 5941(25.48%)
Fill in the Blank 1432(7.77%) 2211(9.48%)
Open Ended Writing 798(4.33%) 1205(5.17%)
Others 100(0.54%) -
Question with Images 7489(40.63%) 9042(38.77%)
 - Single Image 7265(39.42%) 8767(37.59%)
 - Choices within Image 1179(6.40%) 1181(5.06%)
 - Multiple Images 224(1.22%) 275(1.18%)
Question with Explanations 10565(57.33%) 13186(56.54%)
Question with Knowledge 9048(49.09%) 12919(55.40%)
Table 8: The statistic overview of Multi.
A.1 Data Distribution on Question Types
Our benchmark showcases a remarkable diversity in the choice setting of multiple-choice questions, encompassing options that range from 2 to as many as 13. Furthermore, it includes questions that vary in the number of correct answers, from questions with a unique correct option to those with multiple correct choices. We provide the distribution of choices in multiple-choice questions as shown in Table 9. Each row corresponds to a different total number of options available in the questions. The columns represent the frequency of each specific choice option. The table showcases a well-balanced distribution of choices. Notably, the distribution reveals a higher frequency of questions with four choices and a single correct answer, indicating a common format in multiple-choice questions.
Type # choices # A # B # C # D # E,F,G…
SA 2 1819 1376 0 0 0
3 272 287 262 0 0
4 2193 2638 2708 2379 0
5 0 2 7 9 0
MA 3-13 1467 1568 1510 1303 91
Total 2-13 5751 5871 4487 3691 91
Table 9: The choice distribution for multiple-choice questions.
In addition to multiple-choice questions, our benchmark also includes a substantial number of fill-in-the-blank and open-ended questions, creating a diverse and comprehensive range of testing scenarios. Moreover, we have incorporated unique open-response questions that require creative answers, such as drawings. It is important to note that these open-response questions are not included in our formal evaluation and scoring procedures; they are primarily proposed for qualitative research and development in the field of MLLMs. Our benchmark is carefully designed to thoroughly assess and enhance the ability of MLLMs to process and respond to various question types, resembling real-world scenarios.
Appendix B Models
The model specifications are listed in Table 10.
Creator Model # Paras Form Modality Lang Version
FDU MOSS Sun et al. (2023b) 16B Weight T zh, en moss-moon-003-sft
SJTU&AISpeech DFM-2.0 Chen et al. (2022) 70B Weight T zh, en dfm-2.0-70b-preview
LinkSoul-AI Chinese-LLaVA LinkSoul-AI (2023) 7B Weight One zh, en Chinese-LLaVA-Cllama2
THU VisualGLM Du et al. (2022) 6B Weight SI zh, en visualglm-6b
ModelBest VisCPM Hu et al. (2023) 10B Weight SI zh, en VisCPM-Chat
Alibaba Qwen-VL Bai et al. (2023a) 7B Weight MI zh, en Qwen-VL-Chat
OpenGVLab InternVL Chen et al. (2023) 19B Weight One zh, en InternVL-Chat-Chinese-V1.1
01-ai Yi-VL 01.ai (2023) 34B Weight One zh, en Yi-34B-Chat
Google Gemini Team (2023) - API T ML gemini-pro
Google Gemini Vision Team (2023) - API MI ML gemini-pro-vision
OpenAI ChatGPT OpenAI (2022) - API T ML gpt-3.5-turbo-1106
OpenAI GPT-4 OpenAI (2023a) - API T ML gpt-4-1106-preview
OpenAI GPT-4V OpenAI (2023b) - API MI ML gpt-4-vision-preview
Table 10: The list of models evaluated on Multi. We report Modality as how many images can the model take in one turn. Note that those MLLMs commonly support multiple-image input with chatting in several turns. W: accessible through weight. T: pure text LLM, One: only one image in the beginning, SI: single image in each turn, MI: multiple images in one turn. The underline means the model must have an image as input. ML: Multi-lingual.
Appendix C Prompts
The complete collection of prompts designed for evaluation on Multi is shown in Figure 5. One of the prompt pieces in each row is selected according to the evaluation setting and data format. Please note that some prompts will not take effect under certain cases, for instance, the prompt related to knowledge will be omitted if the knowledge is not given.
Figure 5: The prompts for evaluation on Multi.
Appendix D Data Selection Algorithm
We mostly pick questions based on its content length $L_{q}$, calculated with function
$L_{q}=$ $\left(a\times\begin{bmatrix}\mathcal{H}(L_{q,\#\text{characters in question}})\\ \mathcal{H}(L_{q,\#\text{characters in answer}})\\ \mathcal{H}(L_{q,\#\text{characters in analysis}})\end{bmatrix}\right.+$
$\left.b\times\begin{bmatrix}\mathcal{H}(L_{q,\#\text{images in question}})\\ \mathcal{H}(L_{q,\#\text{images in answer}})\\ \mathcal{H}(L_{q,\#\text{images in analysis}})\end{bmatrix}\right)^{\top}\begin{bmatrix}1.0\\ 0.1\\ 0.5\end{bmatrix}$
where $q=1,b=1$ are customized weights.
In the formula above, we use a harmonic mean function $\mathcal{H}$ to normalize content length $L_{q,i}$ of each target value $i$ within each knowledge piece $k$ [^6] Note that for those questions without knowledge information, we simply use a ”null” string as a keyword.
$\mathcal{H}(L_{q,i})=\frac{1}{\frac{1}{L_{q,i}}+\frac{1}{\overline{L_{q,i}}}}=\frac{2L_{q,i}\overline{L_{q,i}}}{L_{q,i}^{2}+\overline{L_{q,i}}^{2}}$
where $\overline{L_{q,i}}$ is the arithmetic average of $L_{q,i}$ for all questions with $k$.
Then we pick $N_{k}$ questions within each knowledge piece $k$.
$N_{k}=\lceil\alpha\times\lg(\#\text{questions of }k)\rceil$
where $\alpha=3$ is a customized parameter.
Now we sort $L_{q,k}={L_{q}:q\in k}$ in descendent order.
Then we assign a pick-up probability to select these questions
$Pr[\text{pick up }q]=\begin{cases}1\text{ , for }q\text{ s.t. }L_{q,k}[0]\\ p\text{ , if }q=1\text{ , for }q\text{ of }L_{q,k}[1:m]\\ \quad\quad\quad\quad\quad\quad\quad\text{ or }L_{q,k}[-m:]\\ p\frac{N_{k}-2m}{\#\text{questions of }k}\text{ , otherwise}\end{cases}$
Appendix E Data Process and Annotation
Initially, we extract a total of 2.7 million questions from the internet. Through an algorithmic selection in the preprocessing stage, we narrow this down to 18,000 questions with wide coverage. During the construction, we conduct two rounds of data annotation and three rounds of automatic checking to ensure the granularity and credibility of every question in our set. In the first round of annotation, we filter out and modify questions based on predefined criteria. The second round of data annotation focuses more on semantic analysis and data enhancement. This post-processing stage significantly increases the number of MA questions by 3.22 times, and the total point proportion of non-SA questions rose from 26.0% to 40.1%. We also remove over 800 similar questions.
E.1 Data Pre-process
The raw data range from HTML, photocopy, hand script, and plain text, and we conduct pre-processing to reduce the load of further annotation. We remove most HTML tags indicating irrelevant content of the question such as alignment, color, etc. We reserve tags for underlines (<u> </u>), and we transfer several tagged styles including bold, italic, and tabular data into markdown format. For some complex tables that cannot be well converted, we save them as a screenshot picture after rendering with HTML.
For photocopy and hand script, we adopt OCR tools to convert text content, crop images, and figures, and integrate them into markdown. We further transcript most of the math functions and chemistry structures into  format, with a small portion remaining as images.
E.2 Data Annotation
Figure 6: A screenshot for the main page of the data annotation platform.
We develop an online platform for data annotation stage. The platform consists of text boxes for editing contents and regions for rendering the text to see the final appearance of the data as shown in Figure 6. We employ skilled human annotators and involve them as authors, primarily undergraduate students from top universities in China familiar with exam quizzes and markdown rules, to undertake this comprehensive task covering various aspects from formatting to semantic analysis:
  • Format Level. Tasks at this level involve the removal of residual HTML tags and the conversion of content into markdown format (refer to examples (1) and (3) in Figure 7). This includes transforming complex mathematical and chemical equations, usually in HTML, into . For this purpose, Mathpix  [^7] https://mathpix.com/snipping-tool is utilized for efficiency. The annotators also correct any character-level errors in text and formulas, often resulting from OCR inaccuracies.
  • Content Level. Annotators split the raw content into distinct sub-questions, segregating parts like the question, answer, and analysis (if presented in raw data). We divide the question content into general and specific parts. The general part includes the problem introduction, background information, or instructions applicable across all sub-questions, while the specific part contains details unique to each sub-question. Annotators also standardize typesetting and image placement, ensuring a consistent format across questions of the same type (e.g., for multiple-choice questions with a single image, the format follows problem content(general) + question content(specific) + [MASK] + [IMAGE_1] + choices).
  • Label Level. Annotators evaluate each question’s difficulty and quality. A question is considered of higher quality if it includes comprehensive content, multiple images, or detailed explanations. Difficulty assessment is subjective. These evaluations aid in curating our Multi-Elite dataset. Annotators also verify information like question type, educational level, and related knowledge pieces.
[…]
In Figure 7, we show several examples of complex formation and modification during the data annotation stage. The markdown, , and HTML format code are remained for better format clarity.
Figure 7: Several data annotation examples when constructing Multi.
E.3 Data Post-process
To collect more challenging data for our benchmark, we adopt several data post-process strategies:
  • Formation. During the data preprocessing stage and annotation stage, we format the questions in a render-friendly manner, and meanwhile reduce the similarity to contents that the MLLMs are trained on. During this stage, we assess if there are any omissions or missing elements.
  • Disambiguration. For blank-filling questions containing multiple [MASK]s, we manually modify those with parallel relations into two sub-questions (refer to example (5) in Figure 8) to determine a unique fixed answer for each question.
  • Distillation. This is completed during our annotation process. We reduce assistance information so that the answer must depend on more detailed analysis (refer to example (4) in Figure 8). In this way, we greatly enhance question difficulty.
  • Transformation. We randomly modify the questions such as from single-choice to blank-filling (refer to example (2) in Figure 8), or convert certain kinds of single-choice questions into multiple-choice ones (refer to example (1) and (5) in Figure 8). Lots of single-choice questions have a list of options and the choices are presented as the combination of those options where only one is correct. We transform those questions into multiple-choice questions where the options become new choices and the correct answer corresponds to the combinations. In this way we successfully increase the scale of multiple-choice questions, improving the diversity of the questions.
Figure 8: Several data augmentation examples when constructing Multi.
In Figure 8, we show several examples of complex formation and modification during the data postprocess stage. English translations of Chinese text are shown for better readability.
Appendix F More Examples
In Figure 9, we show more examples for annotated questions. English translations of Chinese text are shown for better readability.
Figure 9: More example of Multi.